OpenStudy (hoybaby):

how do you gind f(g(n)) if f(n)= 3n-1 and g(n) = n^2 +5?

OpenStudy (anonymous):

already solved

OpenStudy (hoybaby):

how so?

OpenStudy (anonymous):

in the function notation f(x) you would replace in the eqaution whatever you put in the grouping symbols. Suppose f(x)=3x-1 often that is just a number: f(2)=3(2)-1=6-1=5

OpenStudy (hoybaby):

i understand that if ifs f(2) i would subistute in but with another equation like g(n)= n^2+5 i am confused

OpenStudy (anonymous):

You posted this same question earlier

OpenStudy (hoybaby):

nobody responded to it taco bell man

OpenStudy (anonymous):

but you can put anthing inside those grouping symbols, for example suppose gain f(x)=3x-1 then f(x=2)=3(x+2)-1 does this make sense so far

OpenStudy (hoybaby):

well with this problem currently i understand to this point \[f(n^2+5)=\]

OpenStudy (hoybaby):

after what i have is where i am lsot because i am unsure what to do next

OpenStudy (anonymous):

in your case, you have f(n)=3n-1 and g(n)=n^2+5 if you want to find f(g(n)) just put g(n) in for the variable in function f for example\[f \left( g(n) \right)=f \left( x^2+5 \right)=3\left( x^2+5 \right)-1\]

OpenStudy (anonymous):

all you have to do from here is distribute the 3 and simplify\[=3x^2+15-1=3x^2+14\]

OpenStudy (anonymous):

sorry i accidentally changed n to x

OpenStudy (hoybaby):

its fine and i think i understand it better now but my next problem is confusing me more i think, can you help?

OpenStudy (hoybaby):

please?

OpenStudy (anonymous):

yes

OpenStudy (hoybaby):

then the next one is g(n)=n+4 and h(n)= 4n +5 but what is confusing me is this segment here. find g(h(10))

OpenStudy (anonymous):

right here is the link to your previous question. http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e7296ed0b8b247045ca52be

OpenStudy (hoybaby):

ooops sorry taco bell but i didnt get a notifaction that it was answered

OpenStudy (anonymous):

The 10 just means evaluate the function that you find at 10. Let's do it one step at a time.

OpenStudy (anonymous):

haha it is no big deal, i was just letting you know and it has a pretty good detailed explanation

OpenStudy (hoybaby):

so in order to work the 10 i must do h(n)=4(10) +5 ? then mixed this sum into the other equation?

OpenStudy (anonymous):

First just find g(h(n))=g(4n+5) \[g(h(n))=\left( 4n+5 \right)+4=4n+9\]now you can find g(h(10))\[g(h(10))=4(10)+9=49\]

OpenStudy (hoybaby):

so i must substite the problem first in the the other equation before putting the 10 in ?

OpenStudy (anonymous):

no, but is one way of doing it. you could sub the 10 into h first then sub that number, 45, into g

OpenStudy (hoybaby):

is the way you showed me the easiest way?

OpenStudy (hoybaby):

ok i am understanding it way more better then i did 30 minutes ago, i am going to list another problem and do it step by step, i would like you to help me by seeing if i am doing it correctly

OpenStudy (anonymous):

if you are going to evaluate the function at more than one domainn value, you would do it the way I showed you; if you are just going to do it once, the second way might be better. i chose the way i showed you because i thought you wanted the practive in finding function of the form f(h(n))

OpenStudy (anonymous):

ok

OpenStudy (hoybaby):

f(x) =2x g(x)= 4x-1 find f(g(-9))

OpenStudy (hoybaby):

i first would do f(4x-1)=2(4x-1) then f(4x-1)= 8x-2

OpenStudy (hoybaby):

then the final step would be f(g(-9))= 8(-9)-1 which equals -74 so therefore f(g(-9)) = -74

OpenStudy (hoybaby):

is that correct?

OpenStudy (anonymous):

yes, correct!

OpenStudy (anonymous):

the other way would be find g(-9) first: g(-9)=4(-9)-1=-36-1=-37 then find f(g(-9))=f(-37)=2(-37)=-74

OpenStudy (hoybaby):

i also see it that way also, now i know 2 ways to figure it out, much better for me

OpenStudy (hoybaby):

well what do i do if the symbol is squared like this

OpenStudy (hoybaby):

g(t)= t^2 =1 h(t)= t-4

OpenStudy (anonymous):

and what are you trying to find?

OpenStudy (hoybaby):

g(h(4))

OpenStudy (hoybaby):

i think this is only problem i need then i can undertstand all problems that wre close to the ones i discussed

OpenStudy (anonymous):

can I assume g(t)=t^2-1 and not t^2=1

OpenStudy (hoybaby):

oops i meant g(t)=t^2+1

OpenStudy (hoybaby):

what is confusing me is the squared, i am unsure of what to do hence the problem i cant think i can do it

OpenStudy (anonymous):

Method 1: h(4)=4-4=0 then g(h(4))=g(0)-0^2+1=0+1=1 Method 2: g(h(t))=g(t-4)=(t-4)^2+1=t^2-8t+16+1=t^2-8t+17 then g(h(4))=4^2-8(4)-15=16-32+17=1

OpenStudy (anonymous):

remember\[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (hoybaby):

but where do you get the 8t from?

OpenStudy (hoybaby):

in method 2

OpenStudy (anonymous):

using the formula i showed t=right after method 2 where a=t and b=4\[(t-4)^2=t^2-2(t)(4)+4^2\]

OpenStudy (anonymous):

-2*t*4=-8t

OpenStudy (hoybaby):

the 2nd method confuses me but method 1 i totally understand

OpenStudy (anonymous):

you can always just use method 1

OpenStudy (hoybaby):

i believe i understand the method now, now i wanna ask another since it has another variable but i feel that i am abusing your help

OpenStudy (anonymous):

i do this to help out

OpenStudy (hoybaby):

then last question i believe i would ask is f(a)=a-3 g9a)=-a-1 find f(g(-2a))

OpenStudy (hoybaby):

acutlly now that i look at it, it fairly easy

OpenStudy (anonymous):

you do it, i'll check your work then

OpenStudy (hoybaby):

i think what i do is g(a)=-2a-1

OpenStudy (hoybaby):

then -2a-1-3 would be-2a-4 so there fore f(g(-2a))= -2a-4

OpenStudy (anonymous):

you get the mando seal of approval!

OpenStudy (hoybaby):

yay but wait i forgot its g(a)=-a^3

OpenStudy (hoybaby):

but wait,

OpenStudy (hoybaby):

the 2 is a negative and -a^3 is a negative so it will be 2a^3 so it will be 8a

OpenStudy (hoybaby):

so then will be the correct answer be f(g(-2a))= 8a-4?

OpenStudy (anonymous):

wait, what is function f

OpenStudy (hoybaby):

funtcion f is f(a)=a-3

OpenStudy (anonymous):

i do think you made a sign error earlier g(-2a)=-(-2a)-1=2a-1 then f(g(-2a))=f(2a-1)=2a-1-3=2a-4

OpenStudy (hoybaby):

yea i put f(g(-2a))=-2a-1 but i mistanekly wrote the g function wrong its acutllay g(a)=-a^3-1. i forgot the third power, but i know the sign will be positive and will that make 8a?

OpenStudy (hoybaby):

?

OpenStudy (anonymous):

g(-2a)=8a^3-1 f(g(-2a))=8a^3-1-3=8a^3-4

OpenStudy (hoybaby):

wait its 2a^3-3-1, what do i do with the third? do i do 2^3 which is 8 and add the a or just stays 2a^3?

OpenStudy (anonymous):

-(-2a)^3=-(-2)^3a^3=-(-8)a^3=8a^3

OpenStudy (anonymous):

both factors the -2 and the a have to be cubed

OpenStudy (hoybaby):

ohhhh ok thanks, going to be going now, YOUR A AMAZING PERSOn, thanks for the help and KEEP IT UP

OpenStudy (anonymous):

later gtg myself